OpenStudy (austinf):

A country's population in 1992 was 70 million. In 1998 it was 76 million. Estimate the population in 2012 using the exponential growth formula. Round your answer to the nearest million.

OpenStudy (aveline):

Start out by writing your equation. y=a(1+r)^x 76=70(1+r)^6 76/70=(1+r)^6 r=0.0138 y=70(1.0138)^x

OpenStudy (aveline):

Now just plug in 20 for x

OpenStudy (austinf):

So, 92.0754? Correct? @Aveline

OpenStudy (aveline):

That's the number of million people. So it would be about 92,075,435 people. Rounded to the nearest million is 92,000,000

OpenStudy (austinf):

Alright. That's what I thought. :) Thankyou!